OpenStudy (anonymous):

Susan is creating a solution. A formula for the solution calls for 7/10 mL of water. Susan wants to make 2/7 less of the solution than the formula creates. How much water should Susan use? Express your answer as a fraction in simplest form. mL

OpenStudy (agl202):

@Tears1213 Need help?

OpenStudy (anonymous):

yes

OpenStudy (agl202):

First find the greatest common denominator of 10 x 7=?

OpenStudy (anonymous):

70

OpenStudy (agl202):

Greatest common denominator is the common number and both 7 and 10 go into in this case it would be 70 right!

OpenStudy (agl202):

So now 7/10 is the same as 49/70 since all you have to do is multiply 7 by both sides and you get 49/70. Same with 2/7 multiply both sides by 10 and you will get?

OpenStudy (anonymous):

so like 2*10=20 and 10*7 =70 am i right im really bad at mah science is my thing

OpenStudy (agl202):

Yes, ur right. so, that will make it 20/70

OpenStudy (agl202):

Exactly! So now grab both of the fractions and subtract them... So \[\frac{ 49 }{ 70} -\frac{ 20 }{ 70 } =?\]

OpenStudy (anonymous):

ok so 49-20= 29 70-70=0 i failed

OpenStudy (agl202):

Ur numerator is correct, But remember that the denominator has to stay the same. So, u should get: \[\frac{ 29 }{ 70 }\]

OpenStudy (anonymous):

ok thxs

OpenStudy (agl202):

Anytime.